OpenStudy (anonymous):

I HAtE equations like this.! Help Pleas Solve: -4n – 26 = -10

OpenStudy (amistre64):

add, then divide ... simple enough

OpenStudy (anonymous):

I know.! I just hate working with negative numbers.! could you help i'll give yhu a medal.? @amistre64

OpenStudy (amistre64):

show me what your first step is. and i will correct you if you go astray

ganeshie8 (ganeshie8):

do this :- flip all signs since u hate negative numbers

ganeshie8 (ganeshie8):

-4n – 26 = -10 is same as, 4n + 26 = 10

OpenStudy (uri):

@seleenababii Do you understand?

OpenStudy (anonymous):

-4n-26 = -10 -4n = -10+26 -4n = 16 n = 16/-4 n = -4

OpenStudy (uri):

We can do it too,we wanted to know if they undertand how to do it @parmesh.kaur

OpenStudy (anonymous):

thank you guys fofr the help i understand it noww(: @uri and @ganeshie8

ganeshie8 (ganeshie8):

glad to hear you understand :D